OpenStudy (anonymous):

For a CoCl2 hydrate, there's .74 grams of water and .88 grams of anhydrate. How do I figure out the formula?

OpenStudy (anonymous):

Divide the mass of water and anhydrate by the molecular masses of water and CoCl2 respectively to get the number of moles. H2O: \[\frac{ 0.74 }{ 18 }= 0.0411\] CoCl2: \[\frac{ 0.88 }{ 129.8 }= 0.006779\]

OpenStudy (anonymous):

Now for the formula of the hydrate, you need to write it as CoCl2.xH2O To find the x: 0.006779 moles of Cocl2 combine with 0.041 moles of water. I mole of CoCl2 will combine with x moles of water? x= \[\frac{ 0.04111 }{ 0.006779}= 6.06\] So the formula will be \[CoCl _{2}. 6H _{2}O\]